OpenStudy (embracetheweebness98):

Solve the following system of equations and show all work. y = −x2 + 4 y = 2x + 1

OpenStudy (asylum15):

I can walk you through this, ok?

OpenStudy (embracetheweebness98):

ok

OpenStudy (asylum15):

Firstly, do you think we could set these equal?

OpenStudy (asylum15):

Ie Y = Y?

OpenStudy (embracetheweebness98):

whats le?

OpenStudy (asylum15):

ie, in other words

OpenStudy (embracetheweebness98):

ok

OpenStudy (asylum15):

y = −x2 + 4 y = 2x + 1 Both Y, set them equal to each other =)

OpenStudy (embracetheweebness98):

-x2 + 4 = 2x +1

OpenStudy (asylum15):

Correct. Now, any idea what we could do?

OpenStudy (embracetheweebness98):

nope

OpenStudy (asylum15):

Ever heard of a quadratic equation?

OpenStudy (embracetheweebness98):

yes, but i don't remember much

OpenStudy (asylum15):

A quadratic takes the form ax2 + bx + c = 0

OpenStudy (asylum15):

Can you see how we could manipulate -x2 + 4 = 2x +1 To reach ax2 + bx + c = 0 ?

OpenStudy (embracetheweebness98):

-x2 + 2x + 5= 0

OpenStudy (asylum15):

Not quite. We can't have a negative x squared.

OpenStudy (embracetheweebness98):

oh okay

OpenStudy (asylum15):

So if we moved it across the equals?

OpenStudy (embracetheweebness98):

idk

OpenStudy (embracetheweebness98):

what would we move?

OpenStudy (asylum15):

Lets get x squared positive, and manipulate to get a ax2 + bx + c = 0 form

OpenStudy (embracetheweebness98):

okay

OpenStudy (asylum15):

What do you get?

OpenStudy (embracetheweebness98):

Am i changing the whole equation?

OpenStudy (asylum15):

-x2 + 4 = 2x +1 => ax2 + bx + c = 0 form

OpenStudy (embracetheweebness98):

2x-x2+5

OpenStudy (asylum15):

Just a moment

OpenStudy (embracetheweebness98):

ok

OpenStudy (asylum15):

Are you still here? :)

OpenStudy (embracetheweebness98):

Yes, was just about to ask you the same question

OpenStudy (asylum15):

Ok, lets show you how this part is done.

OpenStudy (embracetheweebness98):

ok

OpenStudy (asylum15):

-x2 + 4 = 2x +1 1.) Our x squared is negative, so lets make it positive. We do this by bringing it across the equals. = 4 = x^2 + 2x + 1 Cool?

OpenStudy (embracetheweebness98):

okay

OpenStudy (asylum15):

2.) Now we deal with the X. 2x is the only x in the equation so it's fine. It needs to be beside x^2 to satisfy our ax2 + bx + c = 0 form, which it is. 4 = x^2 + 2x + 1

OpenStudy (asylum15):



OpenStudy (embracetheweebness98):

So we basically just moved it back?

OpenStudy (asylum15):

It didn't move at all. We just brought x^2 over. I'm just showing how we didnt need to touch it.

OpenStudy (embracetheweebness98):

Oh i see it

OpenStudy (embracetheweebness98):

You just took away the "=" sign at the beginning

OpenStudy (asylum15):

Are you with me so far?

OpenStudy (embracetheweebness98):

i think so

OpenStudy (asylum15):

A recap. 1.) We have two equations both Y = For this reason, we can set them equal to each other. 2.) We then can see that we have x^2, 2x and constants, for this reason, we can see that we can manipulate ''-x2 + 4 = 2x +1'' into a quadratic in the form ax^2 + bx + c = 0. 3.) We deal with the X^2 first, then X, then constants. 4.) The X^2 cannot be negative, we move it over the equals. 5.) The 2x is fine, and satisfies the form ax^2 + bx + c = 0 as it is. Ok?

OpenStudy (embracetheweebness98):

ok, does it matter whether we picked the "+4" or the "+1"?

OpenStudy (asylum15):

We haven't gotten to the constants just yet.

OpenStudy (asylum15):

4 = x^2 + 2x + 1 This is where we are. Our X^2 is good! Our 2x is good! Now we have a 4 on the left and a 1 on the right.

OpenStudy (embracetheweebness98):

i mean, if we did 1 = x^2 + 2x + 4, would that matter?

OpenStudy (asylum15):

In essence, the final solution will be x^2 + 2x -3 = 0

jhonyy9 (jhonyy9):

y = −x2 + 4 y = 2x + 1 make y = y and will get -x^2 +4 = 2x +1 divide both sides by -1 and will get x^2 -4 = -2x -1 add to both sides 2x+1 and will get x^2 +2x +1 -4 = 0 x^2 +2x -3 = 0 how you like solve it ? using discriminant or by factored ?

OpenStudy (asylum15):

Embrace, can you see how we get x^2 + 2x - 3 = 0?

jhonyy9 (jhonyy9):

nice job - good luck

OpenStudy (embracetheweebness98):

hold on

OpenStudy (embracetheweebness98):

yeah i see it

OpenStudy (asylum15):

So, this is our final solution. It is a quadratic equation. 

OpenStudy (asylum15):

Now, do you know how to solve a quadratic?

OpenStudy (embracetheweebness98):

we graph it?

OpenStudy (asylum15):

In this case, we want to factorise it, to get two solutions.

OpenStudy (asylum15):

But you're not wrong either, we COULD graph it.

OpenStudy (embracetheweebness98):

okay

OpenStudy (embracetheweebness98):

I'll go with factoring

OpenStudy (asylum15):

It would look something like this.

OpenStudy (asylum15):

If we graphed :)

OpenStudy (embracetheweebness98):

ok, i still prefer factoring beacuse i have to explain this in words and i dont wanna have to explain a graph

OpenStudy (asylum15):

You should always factorise unless asked to graph. :)

OpenStudy (embracetheweebness98):

ok

OpenStudy (asylum15):

We want two numbers that multiply to -3 and add to 2.

OpenStudy (embracetheweebness98):

3 and -1

OpenStudy (asylum15):

Correct! If we were to split up the middle just to show: x² - x + 3x - 3 = 0 With me?

OpenStudy (embracetheweebness98):

kind of

OpenStudy (asylum15):

Have you ever factorised a quadratic before? :)

OpenStudy (embracetheweebness98):

yes

OpenStudy (asylum15):

Ok. So you said 3 and -1. This would mean our x^2 + 2x -3 = 0 becomes: x² - x + 3x - 3 = 0

OpenStudy (embracetheweebness98):

The -1 turned into an -x?

OpenStudy (asylum15):

They are equivalent. -1 = -x in this case.

OpenStudy (asylum15):

You said 3 and -1 -x is the equivalent of your -1. 3x is the equivalent of your 3. :)

OpenStudy (embracetheweebness98):

I am lost XD

OpenStudy (asylum15):

\[x^2 + 2x -3 = 0\]

OpenStudy (embracetheweebness98):

I am so sorry, i dont know what step in solving this we're on XD

OpenStudy (asylum15):

It's no problem. Lets just go back to our x^2 + 2x -3 = 0 Ok? :)

OpenStudy (embracetheweebness98):

okay

OpenStudy (asylum15):

In quadratic equations which are trinomial, the following is true: > The product of the roots is the constant term. > The sum of the roots is the coefficient of the middle term.

OpenStudy (asylum15):

Eg. x^2 - 12x + 27 = 0 What would our two roots be?

OpenStudy (asylum15):

We want 2 numbers that multiply to give +27, and add/subtract to give -12.

OpenStudy (embracetheweebness98):

ok